Thursday, 9 am MT, 11 am ET, 15 December 2016

Logistics

CV-331/AOC-317/GB-bsm, 331 Hub 192.33.117.12##6519 Audio 434-817-6286

Agenda

  • 17B
    • CfP Info (Dana, Bryan, Toney, Mark C.)
    • PST (Rick, Dana)
    • External Tools (Dana)
    • Documentation (Dana)
    • SRP Recruitment (Mark C., Dave)
    • TAC Recruitment (Lewis)
  • AOB

Minutes

Attending: BarryClark, RickLively, MarkClaussen, HeidiMedlin, LorantSjouwerman, ToneyMinter, GarethHunt, DanaBalser, LewisBall

  • 17B
    • CfP Info. The NRAO call is almost ready but we still need the VLA pressure plot, the GBO is coming together, and the LBO call is essentially done. Updated info about the joint programs with external facilities has been sent to the various observatories for review.
    • PST. No progress on general testing. Please send any feedback to DanaBalser by 28 December 2016. GarethHunt is working on several Account tickets which do not have to be released for 17B. The ticket to use the helpdesk for udb-admin mail has been released. There are still a few issues with bounced emails but otherwise it seems to be working.
    • External Tools. A few VLA ECT tickets to be tested. GOST should be ready. No changes for the GBT sensitivity calculator.
    • Documentation. No known changes for PST documentation. DanaBalser will check.
    • SRP Recruitment. MarkClaussen reported on the SRP recruitment progress. No action yet on the 6 vacancies for the extragalactic SRPs (AGN, NGA, EGS, HIZ). For the Galactic SRPs (ISM, SFM, ETP, SSP) 2 of 6 vacancies have been filled.
    • TAC Recruitment. LewisBall reported that he will soon send out invitations to the two TAC vacancies (SFM, NGA).

-- DanaBalser - 2016-12-14
Topic revision: r2 - 2016-12-15, DanaBalser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback